Calculate the energy of a photon of radiation with a wavelength of 2.67 x 10-7 m. What part of the electromagnetic spectrum will this radiation fall in?
1
Expert's answer
2021-09-01T03:11:03-0400

Solution:

Energy of a photon:

E = h × ν

or

E = h × c / λ

where

h is Planck's constant (h = 6.626×10−34 J ⋅ s),

c is the speed of light (c = 2.998×108 m ⋅ s−1),

λ is the wavelength of the radiation (λ = 2.67×10−7 m).


Therefore,

E = (6.626×10−34 J ⋅ s × 2.998×108 m ⋅ s−1) / (2.67×10−7 m) = 7.44×10−19 J

E = 7.44×10−19 J

Ultraviolet (λ = 267 nm)


Answer:

E = 7.44×10−19 J,

ultraviolet (UV).